Susan and Betsy bought sport supplies for summer camp. Susan bought six basketballs and one soccer ball for $27. Betsy bought four basketballs and two soccer balls for $22. How much do basketballs cost? How much do soccer balls cost?
Answer by Cromlix(4381)

Hi there,
Make basketballs = x
Make soccer balls = y
6x + y = 27........(1)
4x + 2y = 22.......(2)
Multiply Eq(1) by 2
12x + 2y = 54......(1)
4x + 2y = 22.......(2)
Subtract Eq(2) from Eq(1)
8x = 32
x = 4
Substitute x = 4 into Eq(2)
4x + 2y = 22.......(2)
4(4) + 2y = 22
16 + 2y = 22
2y = 22 - 16
2y = 6
y = 3
Basketballs cost $4
Soccer balls cost $3
Hope this helps :-)